please help me:
chris has scores of 68,65,75, and 78 on his algebra tests. use a compound inequality to find the scores he can make on his final exam to receive a C in the course. the final exam counts as two tests, and a C is received if the final course average is from 70 to 79.
what do i do???

Answer by nerdybill(7384)

chris has scores of 68,65,75, and 78 on his algebra tests. use a compound inequality to find the scores he can make on his final exam to receive a C in the course. the final exam counts as two tests, and a C is received if the final course average is from 70 to 79.
.
Let x = score on final exam
.
Remembering that the final counts as two tests,
the average would be:
[68+65+75+78+x+x]/6
[68+65+75+78+2x]/6
[286+2x]/6
.
To receive a 'C' the minimum score is:
[286+2x]/6 = 70
286+2x = 420
2x = 134
x = 67
.
To receive a 'C' the maximum score is:
[286+2x]/6 = 79
286+2x = 474
2x = 188
x = 94
.
Therefore, to get a 'C' in the class the score for the final range is:
67 <= x <= 94